OpenStudy (anonymous):

a company uses two vans to transport workers from a free parking lot to the workplace between 7:00 and 9:00a.m. one van has 5 more seats than the other. the smaller van makes two trips every morning while the larger one makes only one trip. the two vans can transport 65 people, maximum. Let x be the seats in the small van while y be the seats for the larger van. how many seats does the larger van have?

OpenStudy (anonymous):

y=x+5 x+y=65 x+x+5=65 2x=60 x=30 y=35

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the ans. 25?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

don you want the method?

OpenStudy (anonymous):

now the two vans can transport 71 people , can u help me ?

OpenStudy (anonymous):

x-5+x-5+x=65 x=25

OpenStudy (anonymous):

x is the number of seats in larger van and x-5 is the number of seats in smaller van

OpenStudy (anonymous):

the question is the same it just now has 71 people

OpenStudy (anonymous):

27 seats

OpenStudy (anonymous):

thank u